OpenStudy (goformit100):

Evaluate 1 + (1+2) + (1+2+3) + (1 +2 + 3+ 4) + … upto n terms.

OpenStudy (goformit100):

@ajprincess

OpenStudy (ajprincess):

Am really sorry. I learnt series long ago.Hwever much i try I dnt seem to remember it:(

OpenStudy (goformit100):

Can you tag any expert in my question please

OpenStudy (goformit100):

Who is free now

OpenStudy (ajprincess):

ya sure:) @ganeshie8, @amistre64, @aravindG pls help:)

OpenStudy (amistre64):

\[\sum\sum n\]maybe?

OpenStudy (aravindg):

Notice that the nth term will be sigma n

OpenStudy (anonymous):

You can write the sum as the following: \[\frac{1(1+1)}{2}+\frac{2(2+1)}{2}+\frac{3(3+1)}{2}+\frac{4(4+1)}{2}+\cdots\] The \(n-\)th term would then be \(\dfrac{n(n+1)}{2}\).

OpenStudy (anonymous):

So far I am at \[S=\frac{ 1 }{ 2 }(1(1+1)+2(2+1)13(3+1)+...+n(n+1))\] looks like \[S= \frac{ 1 }{ 2 } \sum_1^n n(n+1)\]

OpenStudy (aravindg):

yep^^

OpenStudy (goformit100):

Can we use @amistre64 's profile pic her. I mean can we use integration as the sum of the series here ?

OpenStudy (amistre64):

0+1 = 1 1+2 = 3 3+3 = 6 6 + 4 = 10 10 + 5 = 15 1 3 6 10 15 2 3 4 5 1 1 1 \[1+2n+\frac12n(n-1)~:~n=0,1,2,...\]

OpenStudy (amistre64):

the series is discrete tho, not continuous

OpenStudy (goformit100):

Ok thank you all of you .

OpenStudy (anonymous):

\[S_n= \frac{ 1 }{ 12 }\left( n(n+1)(2n+1)+3n(n+1) \right)\] Is what I got from evaluating my stuff above.

OpenStudy (anonymous):

I only checked for the first few terms though, there it seems to hold true.

OpenStudy (amistre64):

\[S_n=1+2n+\frac12n(n-1)~:~n=0,1,2,...\] \[S_n=\frac{2+4n+n^2-n}{2}\] \[S_n=\frac{n^2+3n+2}{2}; would have to let n=n+1 to match siths setup\]

OpenStudy (anonymous):

@Spacelimbus, You can probably prove it by induction.

OpenStudy (goformit100):

are you all satisfied with the discussion . I am fully satisfied so closing the question . Thank you sir and madam.